ORDER: PITOCIN 15 MILLIUNITS/MINUTE AVAILABLE: PITOCIN 30 UNITS IN 500ML HOW MANY ML/HR?
Answer #1

ANSWER.

Calculation of amount(ml/hr)of pitocin required.

         Ordered dose = 15 milliunit/minute.

    Formula to calculate volume(ml) of pitocin needed

                Volume(ml) = Desired dose/Dose in hand *Quantity.

                Here

                   Desired dose = 15 milliunit/minute.

                   Dose in hand = 30 unit

          Convert unit to milliunit for calculation.

                  1 unit =1000 milliunit.

             So

                 Dose in hand = 30 * 1000

                                        = 30000 milliunit.

                        Quantity = 500 ml.

    As per formula

                         Volume (ml) = 15/30000 * 500

                                              = 0.0005 * 500

                                              = 0.25 ml /min.

             Amount (ml/hr) of pitocin required = 0.25 * 60 (minutes)

                                                                    = 15 ml/hr.